OpenStudy (anonymous):

Can anyone help me with drafting a parabola for this equation, I need to know where to mark my graph: y=3/2x^2

OpenStudy (anteater):

First thing you might mark would be the vertex. Do you have an idea of how to find the vertex?

OpenStudy (anonymous):

The vertex is (0,0)

OpenStudy (anteater):

Yes :) So then the next thing you could do is to look at a, which is 3/2.

OpenStudy (anonymous):

Ok, then what is my next step?

OpenStudy (anteater):

If you had just this, y = x^2, you would first mark the vertex (0,0). Then since a = 1, you start at the vertex and go over 1 unit and up 1 unit to mark two more points (1,1) and (-1,1). In this case, a = 3/2 so you again would start at the vertex and go over 1 unit and up how many?

OpenStudy (anonymous):

3

OpenStudy (anteater):

a = ?

OpenStudy (anonymous):

a = 1 so I Mark (-1, 1)?

OpenStudy (anteater):

For your equation, a = 3/2 :) so (-1,3/2) and what other point?

OpenStudy (anonymous):

Can you draw it?

OpenStudy (anonymous):

0

OpenStudy (anteater):

Yes, although I'm not that good an artist :)

OpenStudy (anonymous):

No one is worst than me at drawing even when I am right sometimes the lab does not recognize my graph

OpenStudy (anteater):